OpenStudy (anonymous):

The Township of Brock is growing at a rate of 6.5% per annum. How many people are there in Brock Township now, if there will be 15 000 in 4.5 years? So far i have: A=P(1+i)^n 15 000= P (1+0.65)^4.5

hero (hero):

Okay, so what's the problem? Solve for P.....

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just don't know how to go about solving it

OpenStudy (amistre64):

divide off the 1.65^4.5

OpenStudy (anonymous):

and..

OpenStudy (amistre64):

and hit enter in the calculator

OpenStudy (anonymous):

omg duhh thanks

OpenStudy (amistre64):

\[A = P(1+r)^t\] \[\frac{A}{(1+r)^t} = P\]